### v1

# arXiv:2603.02381v1[math.AP] 02 Mar 2026

Abstract.
Inspired by the work of Cossetti and D’Arca [CD25], we show that the general weighted $L^{p}$-Hardy type inequalities [CD25, Theorems 1.1 and 1.2] and the corresponding identities hold for all $1<p<\infty$, thus extending their results beyond the case $p\geq 2$. In addition, we present a general weighted $L^{p}$-Rellich type inequality with a sharp remainder term for quasilinear second order degenerate elliptic differential operators. In particular, even for the classical Laplacian, these identities appear to be new.
Key words and phrases:
Hardy inequality, Rellich inequality, sharp remainders
2020 Mathematics Subject Classification:
26D10, 35J70
This research is funded by the Committee of Science of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP23490970).
1. Introduction
The study of general weighted Hardy type inequalities dates back to the work of Ghoussoub and Moradifam [GM11], where they obtained a necessary and sufficient condition on a pair of positive radial weights $V(|x|)$ and $W(|x|)$ on a ball $B_{R}$ such that the following inequalities hold for all $u\in C_{0}^{\infty}(B_{R})$:
$\int_{B_{R}}V(|x|)|\nabla u|^{2}dx\geq\int_{B_{R}}W(|x|)|u|^{2}dx.$
Their results improved, extended and unified many Hardy inequalities previously known at that time (e.g., [CKN84, BV97, WW03]). Over the next few years, a huge influx of contributions has been made, see, for instance, [GM13, LLZ19, DLT20, LLZ20, DLL22, DLL24b, FLL25].
Another important work that we would like to mention is by Adimurthi and Sekar [AS06], which might have been one of the first papers where the role of the fundamental solution was studied in the context of Hardy inequalities for general elliptic operators. In particular, they established the following inequality:
$\int_{\Omega}|\nabla u|_{A}^{p}dx-\left(\frac{p-1}{p}\right)^{p}\int_{\Omega}\frac{|\nabla E|_{A}^{p}}{E^{p}}|u|^{p}dx\geq 0,$ (1.1)
where $|\xi|_{A}^{2}=\langle A(x)\xi,\xi\rangle$ for $\xi\in\mathbb{R}^{N}$, $A(x)$ is a $N\times N$ symmetric, uniformly positive definite matrix defined on a bounded domain $\Omega\subset\mathbb{R}^{N}$ and $E$ is a solution to
$\begin{cases}-\operatorname{div}\!\left(|\nabla E|_{A}^{\,p-2}\,A\nabla E\right)=\delta_{0}&\text{in }\Omega,\\[5.69046pt] E=0&\text{on }\partial\Omega.\end{cases}$
Here, $0\in\Omega$ and $\delta_{0}$ is a Dirac mass at 0. Shortly after, the approach showed to be not only applicable to both compact and non-compact manifolds [AS06, Section 4.3, 4.4], but also very useful in other settings [FS08, GKY18, KY18, YEN18a, YEN18b, CAZ21, SYZ25, CD25]. For instance, in [FS08], Frank and Seiringer found out that ground state representations can also be used to obtain a general $L^{p}$-Hardy type inequality and its improvement with a sharp constant. To be precise, suppose $\phi$ is a positive weak solution of the weighted $p$-Laplace equation
$-\operatorname{div}\left(V|\nabla\phi|^{p-2}\nabla\phi\right)=W\phi^{p-1},\quad V>0\quad\text{on}\quad\mathbb{R}^{N}.$ (1.2)
Then, for all complex-valued $u$ with compact support, one has
$\int_{\mathbb{R}^{N}}V|\nabla u|^{p}dx\geq\int_{\mathbb{R}^{N}}W|u|^{p}dx,$ (1.3)
where $\int_{\mathbb{R}^{N}}V|\nabla u|^{p}dx$ and $\int_{\mathbb{R}^{N}}W_{+}|u|^{p}dx$ are assumed to be finite. Interestingly, compared to Bessel pairs in [GM11], the condition (1.2) allows also for non-radial weights $V$ and $W$. Under the restriction to $p\geq 2$, due to the convexity inequality, they also obtained an improvement of (1.3) with a sharp constant [FS08, Formula (2.12)].
Continuing from the results of Adimurthi and Sekar, in [AS06, Open problem 4.1], Adimurthi and Sekar asked whether the constant $\left(\frac{p-1}{p}\right)^{p}$ in (1.1) is optimal. This problem was later resolved by Cowan [COW10], whose work initiated a broader and deeper analysis of general Hardy inequalities, including the determination of best constants and possible improvements.
Building on this, Ghoussoub and Moradifam, in their book [GM13, Chapter 10], addressed the possibility of obtaining a similar characterization of Hardy inequalities in terms of Bessel pairs (as in [GM11]), but for more general elliptic operators [GM13, page 156, Open Problem 10]. This question was partially answered in the same book [GM13] and was completely resolved in [RS25]. For the applications of Bessel pairs in non-Euclidean settings, we refer to [LAM18, RGB22, RS25, D’A24, GJR24, NNP25, SYZ25, CD25].
In the very recent work [CD25], Cossetti and D’Arca generalized and unified many of the previous results [RS25, YZ25, AYZ25, SYZ25]. Their approach brings together the supersolution technique employed in [AS06, FS08, CAZ21, SYZ25] and the method introduced by D’Ambrosio [D’A05], thereby unifying these tools within a common framework. However, their proofs relied on the fundamental algebraic identities [CD25, Proposition 2.1 and 2.3], which hold for $p\geq 2$.
In this paper, we demonstrate that the equivalent algebraic identity (2.2), valid for all $1<p<\infty$, allows us to extend the results of [CD25] beyond the case $p\geq 2$ to the entire range $1<p<\infty$. We remark that identity (2.2) has already proved effective in establishing numerous functional inequalities across a wide range of contexts (see, e.g., [DLL22, DFL+23, DGL+24a, CKL+24, KY24, CT25, YZ25, AYZ25, SYZ25, RSY25]).
Apart from $L^{p}$-Hardy type inequalities, we also establish $L^{p}$-Rellich type inequality with a sharp remainder term. First, let us provide some historical background.
In the aforementioned paper [GM11], it was shown that the Bessel pairs (along with some additional conditions) not only give general weighted Hardy inequalities, but also general Hardy-Rellich inequalities of the following form: let $n\geq 4$ and $W(|x|)$ be a radial Bessel potential such that $r\frac{W_{r}(r)}{W(r)}$ decreases to $\lambda$ and if $\lambda\leq n-2$, then for all $u\in C_{0}^{\infty}(B_{R})$, we have
$\int_{B}|\Delta u|^{2}dx\geq\frac{n^{2}(n-4)^{2}}{16}\int_{B}\frac{u^{2}}{|x|^{4}}dx\\ +\left(\frac{n^{2}}{4}+\frac{(n-\lambda-2)^{2}}{4}\right)\beta(W;R)\int_{B}\frac{W(|x|)}{|x|^{2}}u^{2}dx.$ (1.4)
By choosing appropriate Bessel pairs on $(0,R)$, the inequality (1.4) improves in various ways the results of Adimurthi, Grossi and Santra [AGS06] as well as Tertikas and Zographopoulos [TZ07]. In addition, in the same paper [GM11], more general and higher-order versions of (1.4) were obtained that further improve other results from [TZ07].
Although the result (1.4) does give the classical Rellich inequality and its various improvements, here, we are more interested in obtaining a result in the form of Frank and Seiringer (1.3). In [KY17], Kombe and Yener did exactly that: let $V\in C^{2}\left(\mathbb{R}^{N}\right)$ and $W\in L_{loc}^{1}\left(\mathbb{R}^{N}\right)$ be nonnegative functions. Suppose that $\phi\in C^{\infty}\left(\mathbb{R}^{N}\right)$ is a positive function satisfying the differential inequalities
$\Delta\left(V\left|\Delta\phi\right|^{p-2}\Delta\phi\right)\geq W\phi^{p-1}\quad\text{and}\quad-\Delta\phi>0$
a.e. in $\mathbb{R}^{N}$. Then, there holds
$\int_{\mathbb{R}^{N}}V\left|\Delta u\right|^{p}dx\geq\int_{\mathbb{R}^{N}}W|u|^{p}dx.$ (1.5)
for all real-valued $u\in C_{0}^{\infty}(\mathbb{R}^{N})$. Originally, the result was obtained in a more general setting, namely Baouendi-Grushin (see, Example 2.3), but for simplicity, we state it in the standard Euclidean one. Later, by a similar technique, Ruzhansky and Sabitbek [RS25] extended (1.5) to all complex-valued $u\in C_{0}^{2}(\Omega)$:
$\int_{\Omega}V|\Delta|u||^{p}dx\geq\int_{\Omega}W|u|^{p}dx.$ (1.6)
Since $|\Delta u|\geq|\Delta|u||$, the inequality (1.6) actually gives an improvement to (1.5).
In this paper, we give a sharp remainder formula for (1.5) in the setting of general linear second-order differential operators, which include the standard Laplace operator, Baouendi-Grushin, Heisenberg-Greiner and etc.
2. Preliminaries
In this section, we aim to provide the necessary notation and facts concerning the general linear second-order differential operator as well as the $C_{p}$-functional (the key algebraic identity).
Throughout the paper, we use $\mathcal{I}_{k}$ to denote the identity matrix of order $k$, and $|\cdot|$ to denote the standard Euclidean norm. Let $\sigma:=\left(\sigma_{ij}\right)$, $i=1,\ldots,\ell$, $j=1,\ldots,N$ be an $\ell\times N$ matrix with entries $\sigma_{ij}$ and $\frac{\partial}{\partial x_{j}}\sigma_{ij}$ belonging to $C(\mathbb{R}^{N})$. Now define the vector fields $X_{i}$ as
$X_{i}:=\sum^{N}_{j=1}\sigma_{ij}(x)\frac{\partial}{\partial x_{j}},\quad i=1,\ldots,\ell.$
and the horizontal gradient $\nabla_{\mathcal{L}}$ as follows
$\nabla_{\mathcal{L}}:=\sigma\nabla,$
where $\nabla$ is the usual Euclidean gradient in $\mathbb{R}^{N}$. The assumptions above allow us to naturally define the formal adjoint of $X_{i}$, i.e.
$X^{*}_{i}:=-\sum_{j=1}^{N}\frac{\partial}{\partial x_{j}}\left(\sigma_{ij}(x)\ \cdot\ \right)$
along with $\nabla^{*}_{\mathcal{L}}:=\left(X^{*}_{1},\ldots,X^{*}_{\ell}\right)$. Now we can define the horizontal divergence
$\text{div}_{\mathcal{L}}\left(\cdot\right):=-\nabla^{*}_{\mathcal{L}}\cdot=\text{div}\left(\sigma^{T}\ \cdot\ \right)$
and the linear second-order differential operator
[…]
If $1<p<\infty$, we also have the quasilinear second-order degenerate elliptic differential operator defined as
$\mathcal{L}_{p}u:=-\nabla^{*}_{\mathcal{L}}\cdot\left(|\nabla_{\mathcal{L}}u|^{p-2}\nabla_{\mathcal{L}}u\right).$
The product rule for the horizontal gradient, for two scalar functions $f$ and $g$, is given by
$\nabla_{\mathcal{L}}(fg)=\nabla_{\mathcal{L}}(f)g+f\nabla_{\mathcal{L}}(g).$
Furthermore, we note the identity below, which will be used in the proof of the $L^{p}$-Rellich identity.
Proposition 2.1.
Let $1<p<\infty$ and $\Omega\subseteq\mathbb{R}^{N}$. Then, for all complex-valued $u$ on $\Omega$ and non-trivial real-valued $\phi\in C^{2}(\Omega)$, we have
$\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)=\frac{1}{|\phi|^{p-2}\phi}\biggl[\mathcal{L}(|u|^{p})-(p-1)\frac{\mathcal{L}\phi}{\phi}\,|u|^{p}\\ -2(p-1)\frac{\nabla_{\mathcal{L}}(|u|^{p})\cdot\nabla_{\mathcal{L}}\phi}{\phi}+p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\,|u|^{p}\biggr].$ (2.1)
In order to complete this part of the section, we examine some examples of second-order differential operators.
Example 2.2.
If $\sigma=\mathcal{I}_{N}$, then we get the standard gradient $\nabla_{\mathcal{L}}=\nabla$, Laplacian $\mathcal{L}=\Delta$ and the $p$-Laplacian $\mathcal{L}_{p}=\Delta_{p}$.
Example 2.3.
Split $\mathbb{R}^{N}$ in $z=(x,y)\in\mathbb{R}^{m}\times\mathbb{R}^{k}$ with $m,k\geq 1$, $m+k=N$ and $\gamma\geq 0$. Setting the matrix as
$\sigma=\left(\begin{array}[]{cc}\mathcal{I}_{n}&0\\ 0&|x|^{\gamma}\mathcal{I}_{k}\end{array}\right),$
we get the sub-elliptic gradient $\nabla_{\mathcal{L}}=\left(\nabla_{x},|x|^{\gamma}\nabla_{y}\right)$, where $\nabla_{x}$ and $\nabla_{y}$ are the gradients in the variables $x$ and $y$ respectively. The linear operator $\mathcal{L}$ is called the Baouendi-Grushin operator $\mathcal{L}=\Delta_{x}+|x|^{2\gamma}\Delta_{y}$. We note that if $\gamma=0$ and $k=0$, then we recover the Euclidean setting in Example 2.2.
Example 2.4.
Let $z=(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$. Then, for $|z|:=|(x,y)|$ and $\gamma\geq 1$, we set
$\sigma=\begin{pmatrix}\mathcal{I}_{n}&0&2\gamma y|z|^{2\gamma-2}\\ 0&\mathcal{I}_{n}&-2\gamma x|z|^{2\gamma-2}\end{pmatrix}$.
This gives the vector fields:
$X_{i}=\frac{\partial}{\partial x_{i}}+2\gamma y_{i}|z|^{2\gamma-2}\frac{\partial}{\partial t},\qquad Y_{i}=\frac{\partial}{\partial y_{i}}-2\gamma x_{i}|z|^{2\gamma-2}\frac{\partial}{\partial t},\qquad i=1,\ldots,n.$
In the special cases, if $\gamma=1$, then $\mathcal{L}$ coincides with the sub-Laplacian on the Heisenberg group $\mathbb{H}^{n}$. For $\gamma>1$, the second-order differential operator $\mathcal{L}$ becomes the Greiner operator (see, [GRE79]).
Moving to the next part of this section, we introduce the definition of the $C_{p}$-functional as an algebraic identity.
Definition 2.5.
Let $1<p<\infty$. Then, for $\xi,\eta\in\mathbb{C}^{\ell}$, we define
$C_{p}(\xi,\eta):=|\xi|^{p}-|\xi-\eta|^{p}-p|\xi-\eta|^{p-2}\textnormal{Re}(\xi-\eta)\cdot\overline{\eta}\geq 0.$ (2.2)
Here, we note that [CD25, Propositions 2.1 and 2.2] can be described in terms of the $C_{p}$-functional. Moreover, the identity (2.2) contains both [CD25, Propositions 2.1 and 2.2] as it already holds true for all complex-valued $\xi,\eta\in\mathbb{C}$ (i.e., $\ell=1$) and general complex vectors $\xi,\eta\in\mathbb{C}^{\ell}$.
Lemma 2.6 ([CKL+24, Step 3 of Proof of Theorem 1.2]).
Let $p\geq 2$. Then, for $\xi,\eta\in\mathbb{C}^{\ell}$, we have
$C_{p}(\xi,\eta)\geq c_{1}(p)|\eta|^{p},$
where
$c_{1}(p)=\inf_{(s,t)\in\mathbb{R}^{2}\setminus\{(0,0)\}}\frac{\bigl[t^{2}+s^{2}+2s+1\bigr]^{\frac{p}{2}}-1-ps}{\bigl[t^{2}+s^{2}\bigr]^{\frac{p}{2}}}\in(0,1].$ (2.3)
Lemma 2.7 ([CT25, Lemma 2.2]).
Let $1<p<2\leq\ell$. Then, for $\xi,\eta\in\mathbb{C}^{\ell}$, we have
$C_{p}(\xi,\eta)\geq c_{2}(p)\frac{|\eta|^{2}}{\left(|\xi|+|\xi-\eta|\right)^{2-p}},$
where
$c_{2}(p):=\inf_{s^{2}+t^{2}>0}\frac{\left(t^{2}+s^{2}+2s+1\right)^{\frac{p}{2}}-1-ps}{\left(\sqrt{t^{2}+s^{2}+2s+1}+1\right)^{p-2}(t^{2}+s^{2})}\in\left(0,\frac{p(p-1)}{2^{p-1}}\right].$ (2.4)
Lemma 2.8 ([CT25, Lemma 2.3]).
Let $1<p<2\leq\ell$. Then, for $\xi,\eta\in\mathbb{C}^{\ell}$, we have
$C_{p}(\xi,\eta)\leq c_{3}(p)\frac{|\eta|^{2}}{\left(|\xi|+|\xi-\eta|\right)^{2-p}},$
where
$c_{3}(p):=\sup_{s^{2}+t^{2}>0}\frac{\left(t^{2}+s^{2}+2s+1\right)^{\frac{p}{2}}-1-ps}{\left(\sqrt{t^{2}+s^{2}+2s+1}+1\right)^{p-2}(t^{2}+s^{2})}\in\left[\frac{p}{2^{p-1}},+\infty\right).$ (2.5)
The above lemmata are crucial since they allow us to have the following proposition.
Proposition 2.9.
$C_{p}(\xi,\eta)=0$ if and only if $\eta=0$ for $1<p<\infty$ and $\ell\geq 2$.
3. Main results
In this section, using the algebraic identity (2.2), we prove that [CD25, Theorems 1.1 and 1.2] can be extended to any $1<p<\infty$. We begin with the extension of [CD25, Theorem 1.1].
Theorem 3.1.
Let $1<p<\infty$, $\Omega\subseteq\mathbb{R}^{N}$ be an open subset, $Z:\Omega\to\mathbb{R}^{\ell}$, $V(x)\geq 0$ and $W(x)\in L_{loc}^{1}(\Omega)$ such that $V(x)|Z|^{p}\in L_{loc}^{1}(\Omega)$. Let $\lambda>0$ be a fixed positive constant and assume that there exists a non-trivial real-valued $\phi\in C^{1}(\Omega)$ that solves the following equation in the distributional sense in $\Omega$:
$-\textnormal{div}_{\mathcal{L}}\left(V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\left(\nabla_{\mathcal{L}}\phi\cdot Z\right)Z\right)=\lambda W|\phi|^{p-2}\phi.$ (3.1)
Then, for all complex-valued $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}V|\nabla_{\mathcal{L}}u\cdot Z|^{p}dx=\lambda\int_{\Omega}W|u|^{p}dx+\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot Z,V^{\frac{1}{p}}\phi\nabla_{\mathcal{L}}\left(\frac{u}{\phi}\right)\cdot Z\right)dx.$ (3.2)
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.2) for any $c\in\mathbb{C}$ and $p\geq 2$.
Remark 3.2.
The extremizers are “virtual” in the sense that they make the remainder term vanish, while the integrals appearing in the Hardy identity diverge.
The same technique allows us to extend the Hardy type inequality in all relevant directions [CD25, Theorem 1.2].
Theorem 3.3.
Let $1<p<\infty$, $\Omega\subseteq\mathbb{R}^{N}$ be an open subset, $V(x)\geq 0$ and $W(x)$ be two functions in $L_{loc}^{1}(\Omega)$. Let $\lambda>0$ be a fixed positive constant and assume that there exists a non-trivial real-valued $\phi\in C^{1}(\Omega)$ that solves the following equation in the distributional sense in $\Omega$:
$-\textnormal{div}_{\mathcal{L}}\left(V|\nabla_{\mathcal{L}}\phi|^{p-2}\nabla_{\mathcal{L}}\phi\right)=\lambda W|\phi|^{p-2}\phi.$
Then, for all complex-valued $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}V|\nabla_{\mathcal{L}}u|^{p}dx=\lambda\int_{\Omega}W|u|^{p}dx+\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\nabla_{\mathcal{L}}u,V^{\frac{1}{p}}\phi\nabla_{\mathcal{L}}\left(\frac{u}{\phi}\right)\right)dx.$ (3.3)
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.3) for any $c\in\mathbb{C}$.
Remark 3.4.
Since $C_{p}(\cdot,\cdot)\geq 0$, the identities (3.2) and (3.3) yield the following general weighted $L^{p}$-Hardy type inequalities:
$\int_{\Omega}V|\nabla_{\mathcal{L}}u\cdot Z|^{p}dx\geq\lambda\int_{\Omega}W|u|^{p}dx.$
and
$\int_{\Omega}V|\nabla_{\mathcal{L}}u|^{p}dx\geq\lambda\int_{\Omega}W|u|^{p}dx.$
The next result concerns with the $L^{p}$-Rellich type identity.
Theorem 3.5.
Let $1<p<\infty$, $\Omega\subseteq\mathbb{R}^{N}$ be an open subset, $V(x)\geq 0$ and $W$ be two functions in $L^{1}_{loc}(\Omega)$. Let $\lambda$ be a fixed positive constant and assume that there exists a non-trivial real-valued $\phi\in C^{2}(\Omega)$ with $\frac{\mathcal{L}\phi}{\phi}\geq 0$ that solves the following equation in the distributional sense in $\Omega$:
$\mathcal{L}\left(V|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi\right)=\lambda W|\phi|^{p-2}\phi.$
Then, for all complex-valued $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}V|\mathcal{L}u|^{p}\,dx$ $=\lambda\int_{\Omega}W|u|^{p}dx+\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)dx$
$\quad-p(p-1)\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}dx$
$\quad-p\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)dx.$ (3.4)
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.5) for any $c\in\mathbb{C}$.
Remark 3.6.
Likewise, these extremizers are “virtual” since they eliminate the remainder terms at the cost of integral divergence.
Remark 3.7.
One can observe that $|\nabla_{\mathcal{L}}u|\geq|\nabla_{\mathcal{L}}|u||$ a.e in $\Omega$. In fact, we can take a non-trivial $u(x)=A(x)+iB(x)$, where $A=\operatorname{Re}u$ and $B=\operatorname{Im}u$. Then,
$|\nabla_{\mathcal{L}}|u||^{2}=\left|\frac{A\nabla_{\mathcal{L}}A+B\nabla_{\mathcal{L}}B}{|u|}\right|^{2}.$
Since $|u|^{2}=A^{2}+B^{2}$, by the Cauchy-Schwarz inequality, we get that
$|\nabla_{\mathcal{L}}u|\geq|\nabla_{\mathcal{L}}|u||$
a.e. in $\Omega$. We also refer to [RS25, Lemma 2.4] and [RSS19, Theorem 2.1].
Remark 3.8.
Similarly as for Hardy, since $C_{p}(\cdot,\cdot)\geq 0$, $\frac{\mathcal{L}\phi}{\phi}\geq 0$ and $|\nabla_{\mathcal{L}}u|\geq|\nabla_{\mathcal{L}}|u||$ a.e. in $\Omega$, the identity (3.5) naturally leads to the general $L^{p}$-Rellich inequality:
$\int_{\Omega}V|\mathcal{L}u|^{p}dx\geq\lambda\int_{\Omega}W|u|^{p}dx.$
4. Applications
As [CD25, Applications I-III] are derived from [CD25, Theorems 1.1 and 1.2], the validity of those results for all $1<p<\infty$ directly implies that each of the subsequent applications also carries over to the entire range of $p\in(1,\infty)$.
4.1. Extension of Cossetti and D’Arca results.
Let $d:\Omega\to\mathbb{R}$ be a non-negative and continuous function such that $\nabla_{\mathcal{L}}d\neq 0$ a.e. in $\Omega$ with the following properties:
  1. (1)
    There exists $x_{0}\in\Omega$ such that $d\in C^{\infty}(\Omega\setminus\{x_{0}\})$.
  2. (2)
    $r:=\inf_{x\in\Omega}d(x)$ and $R:=\sup_{x\in\Omega}d(x)$, where $R$ can take the value $\infty$.
  3. (3)
    $V(d)\geq 0$ and $W(d)$ are radial weights defined on the interval $I=(r,R)$.
Additionally, we assume [CD25, H.0-H.3]. Under these assumptions, we have the following Corollary as a direct implication of Theorem 3.1 (see, [CD25, Section 3] for more details).
Corollary 4.1.
Let $1<p<\infty$ and $\phi(d)$ be a solution to (3.1). Then, for all complex-valued $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}V(d)\left|\nabla_{\mathcal{L}}u\cdot\frac{\nabla_{\mathcal{L}}d}{\left|\nabla_{\mathcal{L}}d\right|}\right|^{p}dx=\lambda\int_{\Omega}W(d)|u|^{p}\left|\nabla_{\mathcal{L}}d\right|^{p}dx\\ +\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot\frac{\nabla_{\mathcal{L}}d}{\left|\nabla_{\mathcal{L}}d\right|},V^{\frac{1}{p}}\phi(d)\nabla_{\mathcal{L}}\left(\frac{u}{\phi(d)}\right)\cdot\frac{\nabla_{\mathcal{L}}d}{\left|\nabla_{\mathcal{L}}d\right|}\right)dx.$ (4.1)
Moreover, the “virtual” extremizers of (4.1) are of the form $u=c\phi(d)$ with $c\in\mathbb{C}$ and $p\geq 2$.
Remark 4.2.
After dropping the remainder term and restricting $p\geq 2$, in (4.1), we obtain [CD25, Theorem 3.1]:
$\int_{\Omega}V(d)\left|\nabla_{\mathcal{L}}u\cdot\frac{\nabla_{\mathcal{L}}d}{\left|\nabla_{\mathcal{L}}d\right|}\right|^{p}dx\geq\lambda\int_{\Omega}W(d)|u|^{p}\left|\nabla_{\mathcal{L}}d\right|^{p}dx.$ (4.2)
As we said initially, the authors, in [CD25, Section 3], then apply (4.2) to derive both old and new inequalities [CD25, Section 3]. The identity (4.1) provides an explicit sharp remainder term and extends all of the subsequent results in [CD25, Section 3] to the entire range of $1<p<\infty$. In fact, the same applies for [CD25, Section 4-5].
4.2. Special cases
Here, we discuss some interesting special cases of Theorem 3.3 that are not present in [CD25]. For example, let us consider an open bounded set $\Omega$ of $\mathbb{R}^{N}$ and the following eigenvalue problem:
$\begin{cases}-\mathcal{L}_{p}u=\lambda|u|^{p-2}u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$ (4.3)
where any non-trivial weak solution to (4.3) is called an eigenfunction of the degenerate Dirichlet $p$-Laplacian $\mathcal{L}_{p}u=-\nabla^{*}_{\mathcal{L}}\cdot\left(|\nabla_{\mathcal{L}}u|^{p-2}\nabla_{\mathcal{L}}u\right)$ with a corresponding eigenvalue $\lambda\in\mathbb{R}$. We can also define the first eigenvalue as
$\lambda_{1}(p,\Omega,\mathcal{L})=\inf\left\{\frac{\int_{\Omega}|\nabla_{\mathcal{L}}u|^{p}dx}{\int_{\Omega}|u|^{p}dx}:u\in W_{0}^{1,p}(\Omega)\setminus\{0\}\right\}.$ (4.4)
and the associated eigenfunction we denote as $u_{1}$. Thus, having this in mind, choosing
$V=W=1,\quad\lambda=\lambda_{1}(p,\Omega,\mathcal{L})\quad\text{and}\quad\phi=u_{1}$
in Theorem 3.3, we get the following general identity:
Corollary 4.3.
Let $1<p<\infty$ and $\Omega$ be an open bounded subset of $\mathbb{R}^{N}$. Then, for all complex-valued $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}|\nabla_{\mathcal{L}}u|^{p}dx=\lambda_{1}(p,\Omega,\mathcal{L})\int_{\Omega}|u|^{p}+\int_{\Omega}C_{p}\left(\nabla_{\mathcal{L}}u,u_{1}\nabla_{\mathcal{L}}\left(\frac{u}{u_{1}}\right)\right)dx.$ (4.5)
Moreover, the “virtual” extremizers of (4.5) are of the form $u=cu_{1}$ for any $c\in\mathbb{C}$.
Remark 4.4.
In the special case, identity (4.5) recovers known $L^{p}$-Poincaré identities such as those in [OS20, SY23, AYZ25] for the standard Euclidean and general Baouendi-Grushin cases.
We also obtain some new identities for the $L^{p}$-Rellich inequality. In particular, by taking
$\sigma=\mathcal{I}_{N},\quad V=1\quad\text{and}\quad\phi=|x|^{-\alpha},\quad\text{where}\quad\alpha=\frac{N-2p}{p},$
in (3.5), we obtain the following sharp remainder term of the classical $L^{p}$-Rellich inequality:
Corollary 4.5.
Let $1<p\leq\frac{N}{2}$. Then, for all complex-valued $u\in C_{0}^{\infty}(\mathbb{R}^{N}\backslash\{0\})$, we have
$\int_{\mathbb{R}^{N}}|\Delta u|^{p}dx=C^{p}_{N,p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{2p}}dx+\int_{\mathbb{R}^{N}}C_{p}\left(\Delta u,\Delta u+\frac{C_{N,p}}{|x|^{2}}\right)dx\\ +p(p-1)C^{p-1}_{N,p}\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}}{|x|^{2(p-1)}}\left|\nabla|u|+\frac{N-2p}{p}|u|\frac{x}{|x|^{2}}\right|^{2}dx\\ +pC^{p-1}_{N,p}\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}}{|x|^{2(p-1)}}\left(|\nabla u|^{2}-|\nabla|u||^{2}\right)dx,$ (4.6)
where $C_{N,p}=\frac{N(p-1)(N-2p)}{p^{2}}$. Moreover, the functions $u=c|x|^{-\frac{N-2p}{p}}$ are “virtual” extremizers of (4.6) for all $c\in\mathbb{C}$.
Remark 4.6.
After dropping the remainder terms, we obtain the $L^{p}$-Rellich inequality with a sharp constant:
$\int_{\mathbb{R}^{N}}|\Delta u|^{p}dx\geq\left(\frac{N(p-1)(N-2p)}{p^{2}}\right)^{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{2p}}dx.$
The case $p=2$ of (4.6) also seems to be new.
Corollary 4.7.
Let $N\geq 4$. Then, for all complex-valued $u\in C_{0}^{\infty}(\mathbb{R}^{N}\backslash\{0\})$, we have
$\int_{\mathbb{R}^{N}}|\Delta u|^{2}dx=C^{2}_{N}\int_{\mathbb{R}^{N}}\frac{|u|^{2}}{|x|^{4}}dx+\int_{\mathbb{R}^{N}}\left|\Delta u+\frac{C_{N}}{|x|^{2}}\right|^{2}dx\\ +2C_{N}\int_{\mathbb{R}^{N}}\frac{1}{|x|^{2}}\left|\nabla|u|+\frac{N-4}{4}|u|\frac{x}{|x|^{2}}\right|^{2}dx\\ +2C_{N}\int_{\mathbb{R}^{N}}\frac{|\nabla u|^{2}-|\nabla|u||^{2}}{|x|^{2}}dx,$ (4.7)
where $C_{N}=\frac{N(N-4)}{4}$. Moreover, the functions $u=c|x|^{-\frac{N-4}{2}}$ are “virtual” extremizers of (4.7) for all $c\in\mathbb{C}$.
Remark 4.8.
In the same spirit, we obtain the classical $L^{2}$-Rellich inequality as a consequence of (4.7):
$\int_{\mathbb{R}^{N}}|\Delta u|^{2}dx\geq\left(\frac{N(N-4)}{4}\right)^{2}\int_{\mathbb{R}^{N}}\frac{|u|^{2}}{|x|^{4}}dx.$
Remark 4.9.
The above results are almost identical in structure to the general case for the $\mathcal{L}$ operator. However, we state the Euclidean ones for simplicity and to emphasize the novelty even for the standard Laplacian $\Delta$.
5. Proofs of Theorems 3.1 and 3.3
This section is devoted to the proofs of Theorems 3.1 and $3.$
Proof of Theorem 3.1.
We know that
$C_{p}(\xi,\eta)=|\xi|^{p}-|\xi-\eta|^{p}-p|\xi-\eta|^{p-2}\text{Re}\left(\xi-\eta\right)\cdot\overline{\eta}.$
for $1<p<\infty$ and $\xi,\eta\in\mathbb{C}^{n}$. Let $\xi:=f$ and $\eta:=f-g$. Then
$C_{p}(f,f-g)=|f|^{p}+(p-1)|g|^{p}-p\text{Re}|g|^{p-2}g\overline{f}.$ (5.1)
Now take $\Omega\subseteq\mathbb{R}^{N}$, complex-valued $u\in C_{0}^{\infty}(\Omega)$, $V(x)\geq 0$, $W(x)\in L_{loc}^{1}(\Omega)$, $Z:\Omega\to\mathbb{R}^{\ell}$ and a non-trivial real-valued $\phi\in C^{1}(\Omega)$. Define
$f:=V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot Z\in L^{p}\left(\Omega\right)\quad\text{and}\quad g:=\frac{\nabla_{\mathcal{L}}\phi\cdot Z}{\phi}uV^{\frac{1}{p}}\in L^{p}(\Omega).$ (5.2)
By direct computation, we get
$p\text{Re}\int_{\Omega}|g|^{p-2}g\overline{f}dx$ $=p\text{Re}\int_{\Omega}\frac{V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\nabla_{\mathcal{L}}\phi\cdot Z}{|\phi|^{p-2}\phi}|u|^{p-2}u\overline{\nabla_{\mathcal{L}}u}\cdot Zdx$
$=\int_{\Omega}\nabla_{\mathcal{L}}|u|^{p}\cdot\frac{V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\nabla_{\mathcal{L}}\phi\cdot Z}{|\phi|^{p-2}\phi}Zdx$
$=\int_{\Omega}\nabla_{\mathcal{L}}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)\cdot\left(V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\nabla_{\mathcal{L}}\phi\cdot Z\right)Zdx$
$\quad+(p-1)\int_{\Omega}|u|^{p}\frac{|\nabla_{\mathcal{L}}\phi\cdot Z|^{p}}{|\phi|^{p}}Vdx.$ (5.3)
Let us assume that there exists a non-trivial real-valued $\phi\in C^{1}(\Omega)$ that solves the following equation in the distributional sense in $\Omega$:
$-\text{div}_{\mathcal{L}}\left(V|\nabla_{\mathcal{L}}\phi\cdot Z|\left(\nabla_{\mathcal{L}}\phi\cdot Z\right)Z\right)=\lambda W|\phi|^{p-2}\phi,$
which is equivalent to
$\int_{\Omega}V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\left(\nabla_{\mathcal{L}}\phi\cdot Z\right)\left(\nabla_{\mathcal{L}}\varphi\cdot Z\right)dx=\lambda\int_{\Omega}W|\phi|^{p-2}\phi\ \varphi dx$ (5.4)
for all $\varphi\in C^{\infty}_{c}(\Omega)$. Substituting $\varphi=\frac{|u|^{p}}{|\phi|^{p-2}\phi}$ to (5.4), we have that
$\int_{\Omega}V|\nabla_{\mathcal{L}}\phi\cdot Z|^{p-2}\left(\nabla_{\mathcal{L}}\phi\cdot Z\right)\left(\nabla_{\mathcal{L}}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)\cdot Z\right)dx=\lambda\int_{\Omega}W|u|^{p}dx.$ (5.5)
Using (5.5) in (5.3), we get
$p\text{Re}\int_{\Omega}|g|^{p-2}g\overline{f}dx=\lambda\int_{\Omega}W|u|^{p}dx+(p-1)\int_{\Omega}|u|^{p}\frac{|\nabla_{\mathcal{L}}\phi\cdot Z|^{p}}{|\phi|^{p}}Vdx.$ (5.6)
Now let us integrate (5.1) over $\Omega$:
$\int_{\Omega}C_{p}(f,f-g)dx=\int_{\Omega}|f|^{p}dx+(p-1)\int_{\Omega}|g|^{p}dx-p\text{Re}\int_{\Omega}|g|^{p-2}g\overline{f}dx.$ (5.7)
Putting (5.2) and (5.6) to (5.7), we obtain
$\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot Z,V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot Z-\frac{\nabla_{\mathcal{L}}\phi\cdot Z}{\phi}uV^{\frac{1}{p}}\right)dx=\int_{\Omega}V|\nabla_{\mathcal{L}}u\cdot Z|^{p}dx\\ +(p-1)\int_{\Omega}|u|^{p}\frac{|\nabla_{\mathcal{L}}\phi\cdot Z|^{p}}{|\phi|^{p}}Vdx-\lambda\int_{\Omega}W|u|^{p}dx-(p-1)\int_{\Omega}|u|^{p}\frac{|\nabla_{\mathcal{L}}\phi\cdot Z|^{p}}{|\phi|^{p}}Vdx\\ =\int_{\Omega}V|\nabla_{\mathcal{L}}u\cdot Z|^{p}dx-\lambda\int_{\Omega}W|u|^{p}dx.$
Dropping the remainder term gives us
$\int_{\Omega}V|\nabla_{\mathcal{L}}u\cdot Z|^{p}dx\geq\lambda\int_{\Omega}W|u|^{p}dx.$ (5.8)
Finally, since
$C_{p}(\xi,\eta)=0\iff\eta=f-g=0$
and
$f-g=V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\cdot Z-\frac{\nabla_{\mathcal{L}}\phi\cdot Z}{\phi}uV^{\frac{1}{p}}=V^{\frac{1}{p}}\phi\nabla_{\mathcal{L}}\left(\frac{u}{\phi}\right)\cdot Z=0,$
we have that the remainder term vanishes (or that the inequality (5.8) achieves equality) if and only if $u=c\phi$ for any $c\in\mathbb{C}$. ∎
Proof of Theorem 3.3.
Let $u$ and $V,\phi$ be as in the assumptions of Theorem 3.3. Taking
$f:=V^{\frac{1}{p}}\nabla_{\mathcal{L}}u\in L^{p}\left(\Omega\right)\quad\text{and}\quad g:=\frac{\nabla_{\mathcal{L}}\phi}{\phi}uV^{\frac{1}{p}}\in L^{p}(\Omega)$ (5.9)
and proceeding in the same way as in the proof of Theorem 3.1, we get
$\int_{\Omega}C_{p}\!\left(V^{\frac{1}{p}}\nabla_{\mathcal{L}}u,V^{\frac{1}{p}}\nabla_{\mathcal{L}}u-\frac{\nabla_{\mathcal{L}}\phi}{\phi}uV^{\frac{1}{p}}\right)dx$ $=\int_{\Omega}V|\nabla_{\mathcal{L}}u|^{p}dx-\lambda\int_{\Omega}W|u|^{p}dx.$
Since $C_{p}\geq 0$, we get the inequality
$\int_{\Omega}V|\nabla_{\mathcal{L}}u|^{p}dx\geq\lambda\int_{\Omega}W|u|^{p}dx.$ (5.10)
Finally, equality in (5.10) holds iff
$V^{\frac{1}{p}}\nabla_{\mathcal{L}}u-\frac{\nabla_{\mathcal{L}}\phi}{\phi}uV^{\frac{1}{p}}=0.$
That is,
$V^{\frac{1}{p}}\phi\,\nabla_{\mathcal{L}}\!\left(\frac{u}{\phi}\right)=0,$
which gives $u=c\phi$ for some constant $c\in\mathbb{C}$. ∎
6. Proofs of Proposition 2.1, Theorem 3.5 and Corollary 4.5
First, let us provide the proof of Proposition 2.1. Then, we will proceed with Theorem 3.5.
Proof of Proposition 2.1.
We are given that $1<p<\infty$, $\Omega\subseteq\mathbb{R}^{N}$. Now let us take a complex-valued $u$ on $\Omega$ and a non-trivial real-valued $\phi\in C^{2}(\Omega)$. Then, by direct calculations, we get
$\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)=\mathcal{L}\left(v\Phi\right),$
where
$v=|u|^{p}\quad\text{and}\quad\Phi=\frac{1}{|\phi|^{p-2}\phi}.$
Then, by the product rule, we get
$\mathcal{L}(v\Phi)=\mathcal{L}(v)\Phi+v\mathcal{L}(\Phi)+2\nabla_{\mathcal{L}}v\cdot\nabla_{\mathcal{L}}\Phi.$
Let us calculate each part term by term. We start with $\mathcal{L}(v)\Phi$:
$\mathcal{L}(v)\Phi=\mathcal{L}(|u|^{p})\Phi.$
To evaluate the rest, we note that
$|\phi|^{p}\Phi=\phi.$
Therefore, we have
$\mathcal{L}\left(|\phi|^{p}\Phi\right)$ $=\mathcal{L}(|\phi|^{p})\Phi+|\phi|^{p}\mathcal{L}(\Phi)+2\nabla_{\mathcal{L}}|\phi|^{p}\cdot\nabla_{\mathcal{L}}\Phi$
$=\left[p(p-1)|\phi|^{p-2}|\nabla_{\mathcal{L}}\phi|^{2}+p|\phi|^{p-2}\phi\mathcal{L}\phi\right]\Phi+|\phi|^{p}\mathcal{L}(\Phi)+2\nabla_{\mathcal{L}}|\phi|^{p}\cdot\nabla_{\mathcal{L}}\Phi$
$=p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi}+p\mathcal{L}\phi+|\phi|^{p}\mathcal{L}(\Phi)+2\nabla_{\mathcal{L}}|\phi|^{p}\cdot\nabla_{\mathcal{L}}\Phi$
$=p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi}+p\mathcal{L}\phi+|\phi|^{p}\mathcal{L}(\Phi)$
$\quad+2\left(p|\phi|^{p-2}\phi\nabla_{\mathcal{L}}\phi\right)\cdot\left((1-p)|\phi|^{-p}\nabla_{\mathcal{L}}\phi\right)$
$=-p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi}+p\mathcal{L}\phi+|\phi|^{p}\mathcal{L}\Phi=\mathcal{L}\phi.$
Since $\phi\neq 0$, this gives that
$\mathcal{L}\Phi=\Phi\left[-(p-1)\frac{\mathcal{L}\phi}{\phi}+(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\right]$
Combining all pieces together, we get
$\mathcal{L}\left(v\Phi\right)$ $=\mathcal{L}(|u|^{p})\Phi+|u|^{p}\Phi\left[-(p-1)\frac{\mathcal{L}\phi}{\phi}+p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\right]$
$\quad-2(p-1)\Phi\frac{\nabla_{\mathcal{L}}(|u|^{p})\cdot\nabla_{\mathcal{L}}\phi}{\phi},$
which after factoring $\Phi$, gives us (2.1). ∎
Proof of Theorem 3.5.
Let us recall the identity (2.1):
$\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)=\frac{1}{|\phi|^{p-2}\phi}\biggl[\mathcal{L}(|u|^{p})-(p-1)\frac{\mathcal{L}\phi}{\phi}\,|u|^{p}\\ -2(p-1)\frac{\nabla_{\mathcal{L}}(|u|^{p})\cdot\nabla_{\mathcal{L}}\phi}{\phi}+p(p-1)\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\,|u|^{p}\biggr].$
Multiplying by $|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi$, gives us
$\mathcal{L}\!\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)\,|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi$
$\quad=\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\mathcal{L}(|u|^{p})-(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{\mathcal{L}\phi}{\phi}\,|u|^{p}$
$\quad\quad-2(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{\nabla_{\mathcal{L}}(|u|^{p})\cdot\nabla_{\mathcal{L}}\phi}{\phi}+p(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\,|u|^{p}.$
Thus, we have
$\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\mathcal{L}(|u|^{p})$
$\quad=\mathcal{L}\!\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)\,|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi+(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{\mathcal{L}\phi}{\phi}\,|u|^{p}$
$\quad\quad+2(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{\nabla_{\mathcal{L}}(|u|^{p})\cdot\nabla_{\mathcal{L}}\phi}{\phi}-p(p-1)\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}\,|u|^{p}.$ (6.1)
On the other hand, similarly as in the proofs of Theorem 3.1 and 3.3, by direct calculation, we get
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}+(p-1)\,V\frac{|\mathcal{L}\phi|^{p}}{|\phi|^{p}}\,|u|^{p}-p\operatorname{Re}\!\left(V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\,|u|^{p-2}u\,\overline{\mathcal{L}u}\right).$ (6.2)
Using the identity
$-p|u|^{p-2}\operatorname{Re}(u\overline{\mathcal{L}u})=-\mathcal{L}(|u|^{p})+p|u|^{p-2}|\nabla_{\mathcal{L}}u|^{2}+p(p-2)|u|^{p-4}|\operatorname{Re}(u\overline{\nabla_{\mathcal{L}}u})|^{2}$
in (6), we get
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}+(p-1)\,V\frac{|\mathcal{L}\phi|^{p}}{|\phi|^{p}}\,|u|^{p}-V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\mathcal{L}(|u|^{p})$
$\quad\quad+pV\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}|\nabla_{\mathcal{L}}u|^{2}+p(p-2)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-4}\left|\text{Re}\left(u\overline{\nabla_{\mathcal{L}}u}\right)\right|^{2}.$ (6.3)
Putting (6) in (6), we obtain
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}+(p-1)\,V\frac{|\mathcal{L}\phi|^{p}}{|\phi|^{p}}\,|u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi$
$\quad\quad-(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\frac{\mathcal{L}\phi}{\phi}|u|^{p}-2(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\frac{\nabla_{\mathcal{L}}\left(|u|^{p}\right)\cdot\nabla_{\mathcal{L}}\phi}{\phi}$
$\quad\quad+p(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}|u|^{p}+pV\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}|\nabla_{\mathcal{L}}u|^{2}$
$\quad\quad+p(p-2)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-4}\left|\text{Re}\left(u\overline{\nabla_{\mathcal{L}}u}\right)\right|^{2}.$
Simplifying:
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi-2(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\frac{\nabla_{\mathcal{L}}\left(|u|^{p}\right)\cdot\nabla_{\mathcal{L}}\phi}{\phi}$
$\quad\quad+p(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}|u|^{p}+pV\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}|\nabla_{\mathcal{L}}u|^{2}$
$\quad\quad+p(p-2)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-4}\left|\text{Re}\left(u\overline{\nabla_{\mathcal{L}}u}\right)\right|^{2}.$
Let $G:=V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}$. Since
$\text{Re}\left(u\overline{\nabla_{\mathcal{L}}u}\right)=|u|\nabla_{\mathcal{L}}(|u|)\quad\text{and}\quad\nabla_{\mathcal{L}}|u|^{p}=p|u|^{p-2}\text{Re}\left(u\overline{\nabla_{\mathcal{L}}u}\right)=p|u|^{p-2}|u|\nabla_{\mathcal{L}}(|u|),$
we get
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi+p(p-1)G\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}|u|^{p}$
$\quad\quad-2p(p-1)G|u|^{p-2}\left(\nabla_{\mathcal{L}}|u|\cdot\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right)+pG|u|^{p-2}|\nabla_{\mathcal{L}}u|^{2}$
$\quad\quad+p(p-2)G|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|\right|^{2}$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi+p(p-1)G\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}|u|^{p}$
$\quad\quad-2p(p-1)G|u|^{p-2}\left(\nabla_{\mathcal{L}}|u|\cdot\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right)+pG|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|\right|^{2}$
$\quad\quad+pG|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)+p(p-2)G|u|^{p-2}|\nabla_{\mathcal{L}}|u||^{2}$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi+p(p-1)G\frac{|\nabla_{\mathcal{L}}\phi|^{2}}{\phi^{2}}|u|^{p}$
$\quad\quad-2p(p-1)G|u|^{p-2}\left(\nabla_{\mathcal{L}}|u|\cdot\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right)+p(p-1)G|u|^{p-2}|\nabla_{\mathcal{L}}|u||^{2}$
$\quad\quad+pG|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi$
$\quad\quad+p(p-1)G|u|^{p-2}\left[|\nabla_{\mathcal{L}}|u||^{2}-2\nabla_{\mathcal{L}}|u|\cdot\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi+\frac{|u|^{2}}{\phi^{2}}|\nabla_{\mathcal{L}}\phi|^{2}\right]$
$\quad\quad+pG|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi+p(p-1)G|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}$
$\quad\quad+pG|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right).$
Making the change of variables $G=V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}$, we have
$C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)$
$\quad=V|\mathcal{L}u|^{p}-V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi$
$\quad\quad+p(p-1)V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}$
$\quad\quad+pV\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right).$
Integrating over $\Omega$, we obtain
$\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)dx$
$\quad=\int_{\Omega}V|\mathcal{L}u|^{p}\,dx-\int_{\Omega}V\mathcal{L}\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right)|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi dx$
$\quad\quad+p(p-1)\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}dx$
$\quad\quad+p\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)dx.$
Applying integration by parts and the fact that there exists a non-trivial real-valued $\phi\in C^{2}\left(\Omega\right)$ that solves the following equation in the distributional sense in $\Omega$:
$\mathcal{L}\left(V|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi\right)=\lambda W|\phi|^{p-2}\phi,$
we obtain
$\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)dx$
$\quad=\int_{\Omega}V|\mathcal{L}u|^{p}dx-\lambda\int_{\Omega}W|u|^{p}dx$
$\quad\quad+p(p-1)\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}dx$
$\quad\quad+p\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)dx.$
Rewriting
$\int_{\Omega}V|\mathcal{L}u|^{p}dx$
$\quad=\lambda\int_{\Omega}W|u|^{p}dx+\int_{\Omega}C_{p}\left(V^{\frac{1}{p}}\mathcal{L}u,V^{\frac{1}{p}}\mathcal{L}u-V^{\frac{1}{p}}\frac{\mathcal{L}\phi}{\phi}u\right)dx$
$\quad\quad-p(p-1)\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left|\nabla_{\mathcal{L}}|u|-\frac{|u|}{\phi}\nabla_{\mathcal{L}}\phi\right|^{2}dx$
$\quad\quad-p\int_{\Omega}V\frac{|\mathcal{L}\phi|^{p-2}\mathcal{L}\phi}{|\phi|^{p-2}\phi}|u|^{p-2}\left(|\nabla_{\mathcal{L}}u|^{2}-|\nabla_{\mathcal{L}}|u||^{2}\right)dx.$
Finally, we set $\frac{\mathcal{L}\phi}{\phi}>0$ and use the fact that $|\nabla_{\mathcal{L}}u|\geq|\nabla_{\mathcal{L}}|u||$ (see, [RS25, Lemma 2.4] and [RSS19, Theorem 2.1]) to complete the proof. ∎
Proof of Corollary 4.5.
We have that $\sigma=\mathcal{I}_{k}$, $V=1$ and $\phi=|x|^{-\alpha}$ with $\alpha=\frac{N-2p}{p}$. By direct calculations, we have
$\phi^{\prime}(r)=-\alpha r^{-\alpha-1},\qquad\phi^{\prime\prime}(r)=\alpha(\alpha+1)r^{-\alpha-2}.$
Hence,
$\Delta\phi=\phi^{\prime\prime}+\frac{n-1}{r}\phi^{\prime}=\alpha(\alpha+1)r^{-\alpha-2}+(n-1)(-\alpha r^{-\alpha-2})=\alpha(\alpha-n+2)r^{-\alpha-2}.$
Let us write this as follows:
$\Delta\phi=c_{1}r^{-\alpha-2},\qquad c_{1}:=\alpha(\alpha-n+2).$
Simplifying, we get
$c_{1}=-A:=\frac{n(p-1)(n-2p)}{p^{2}}>0,\qquad(n>2p).$
Since $\Delta\phi=-Ar^{-\alpha-2}$ and $A>0$,
$|\Delta\phi|^{p-2}\Delta\phi=(Ar^{-\alpha-2})^{p-2}(-Ar^{-\alpha-2})=-A^{p-1}r^{-(\alpha+2)(p-1)}.$
Setting
$g(r):=|\Delta\phi|^{p-2}\Delta\phi=Kr^{-\beta},\qquad K:=-A^{p-1},\qquad\beta:=(\alpha+2)(p-1)=\frac{n(p-1)}{p}.$
The calculations for $\Delta g$ are the same as for $\Delta\phi$:
$\Delta g=\beta(\beta-n+2)Kr^{-\beta-2},$
which means that
$\Delta\bigl(|\Delta\phi|^{p-2}\Delta\phi\bigr)=\beta(\beta-n+2)Kr^{-\beta-2}.$
Computing the constants, we obtain
$\beta(\beta-n+2)K=(-A)(-A^{p-1})=A^{p}$
and thus
$\Delta\bigl(|\Delta\phi|^{p-2}\Delta\phi\bigr)=A^{p}r^{-\beta-2}.$
Since
$-\beta-2=-2p-\alpha(p-1)$
and
$|\phi|^{p-2}\phi=r^{-\alpha(p-1)},$
we finally derive that
$\Delta\bigl(|\Delta\phi|^{p-2}\Delta\phi\bigr)=A^{p}|x|^{-2p}|\phi|^{p-2}\phi.$
Therefore,
$\lambda W=A^{p}|x|^{-2p}.$
Other calculations follow directly:
$\frac{\nabla\phi}{\phi}=-\alpha\frac{x}{|x|^{2}},\quad\frac{\Delta\phi}{\phi}=-A\frac{1}{|x|^{2}},\quad\frac{|\Delta\phi|^{p-2}\Delta\phi}{|\phi|^{p-2}\phi}=-A^{p-1}\frac{1}{|x|^{2(p-1)}}.$
Combining everything together, we get (4.6). ∎